1. Solve the system by the elimination method. (Enter your answers as a comma-separated list. If the system is inconsistent, enter INCONSISTENT. If the system is dependent, enter DEPENDENT.)

3x + y = 19
x + 2y = 13

(x, y)

=

2. Solve the system by graphing. (Enter your answers as a comma-separated list. If the system is inconsistent, enter INCONSISTENT. If the system is dependent, enter DEPENDENT.)

x + y = 9
xy = −9

(x, y)=

3. Solve the system by the elimination method. (Enter your answers as a comma-separated list. If the system is inconsistent, enter INCONSISTENT. If the system is dependent, enter DEPENDENT.)

x + y = 6

2x + 3y = 16

(x, y)=
Answer #1

Q 1)  

Given equations
3x + y = 19 ------ (1)
x + 2y = 13 ------ (2)
Multiplying eq (1) with 2 we get  
2 x ( 3x + y = 19)
6x + 2y = 38 ------- (3)
eq (3) minus eq (2)
6x + 2y = 38 minus x + 2y = 13 we get
5x + 0 = 25
5x = 25
x = 25 / 5 = 5
Substituting x = 5 in eq (1) we get
3 (5) + y = 19
15 +y =19
y = 19 - 15
y = 4

( x , y ) = ( 5 , 4) DEPENDENT

Q 2)

Given equations
x + y = 9 ----- (1)
- x - y = - 9 --- (2)

As eq (2) and eq (1) are similar as eq (1) multiplied by  -1 gives eq (2), therefore we cannot calculate the values of ( x, y) INCONSISTENT

Q 3)

Given equations are
x + y = 6 ---- (1)
2x + 3y = 16 ---- (2)
By multiplying eq (1) by 2 we get
2 x ( x + y = 6)
2x + 2y = 12 ----- (3)
eq (2) minus eq (3) we get
2x + 3y = 16 minus 2x + 2y = 12
0 + y = 4
y = 4
Substituting y=4 in eq (1) we get
x + 4 = 6
x = 6 - 4
x = 2
therefore

( x , y) = ( 2 , 4) DEPENDENT


NOTE -- eq in the above answers represents equation